OpenStudy (anonymous):

A 2.5 × 10^3 kg car requiers 5.00 kJ of work to move from rest to some final velocity. During this time, the car travels 25.0 m. what is the final velocity of the car?

OpenStudy (anonymous):

w= work = force x distance = F d =5 kJ F = mass x acceleration = m a work = change in kinetic energy = (1/2) m v^2 v = sqrt(2 w / m) = sqrt( 10x10^3 N-m/ 2.5 x 10^3 kg) = 2 m/s